What is the difference between mM, uM, and nM when doing PCR

and what do they stand for (ex. uM same as micromolar?)
Answer #1

Ans:- These are three different concentration units mM (milli Molar = 10-3 molars) means any substance for example "A" is present at a concentration of one thousand of a mole per litre.

uM (micro Molar = 10-6) means any substance for example "A" is present at a concentration of one-millionth of a mole per litre.

nM (nano Molar = 10-9) means any substance for example "A" is present at a concentration of one-billionth of a mole per litre.